Determine the force of gravitational attraction between the earth (5.98 x1024


kg) and a 100 kg football player, when he is at a distance of 6.38 x 106 m from


earth’s center.

1
Expert's answer
2021-12-29T12:34:16-0500

Given:

"M_E=\rm 5.98*10^{24}\:\rm kg"

"m=\rm 100\:\rm kg"

"R=\rm 6.38*10^{6}\:\rm kg"

"F-?"


The law of universe gravitation says

"F=G\frac{M_Em}{R^2}\\
=\rm 6.67*10^{-11}\frac{5.98*10^{24}*100}{(6.38*10^6)^2}=980\: N"